1 LBC Corporation makes and sells a product called Product WZ. Each unit of Product WZ requires 4.4 hours of direct labor at the rate of $15.00 per direct labor-hour. Management would like you to prepare a Direct Labor Budget for June. The budgeted direct labor cost per unit of Product WZ would be: Multiple Choice • $66.00 • $9.80 • $15.00 • $42.20

2. Teel Printing uses two measures of activity, press runs and book set-ups, in the cost formulas in its budgets and performance reports. The cost formula for wages and salaries is $4,100 per month plus $405 per press run plus $955 per book set-up. The company expected its activity in July to be 219 press runs and 126 book set-ups, but the actual activity was 216 press runs and 125 book set-ups. The actual cost for wages and salaries in July was $211,780. The spending variance for wages and salaries in July would be closest to: Multiple Choice • $825 U • $1,345 F • $825 F • $1,345 U

3 Courington Detailing's cost formula for its materials and supplies is $2,050 per month plus $11 per vehicle. For the month of August, the company planned for activity of 85 vehicles, but the actual level of activity was 45 vehicles. The actual materials and supplies for the month was $2,880. The spending variance for materials and supplies in August would be closest to: Multiple Choice • $335 F • $105 F • $335 U • $105 U
Answer #1

1) Budgeted direct labor cost per unit = 4.4*15 = 66

So answer is a) $66

2) Spending variance for wages and salaries = (4100+216*405+955*125)-211780 = 825 U

So answer is a) $825 U

3) Spending variance for materials and supplies = (2050*11*45)-2880 = 335 U

So answer is c) $335 U